OpenStudy (anonymous):

Find the distance from (6,2) to the line 2x - 3y = 6 I used this formula: (Ax + By + C) / (sqrt. A^2 + B^2) and got 0 for the numerator. Did I do the problem correctly?

OpenStudy (raden):

you just forgot the absolut sign, it is |(Ax + By + C) / (sqrt. A^2 + B^2)|

OpenStudy (raden):

first, 2x - 3y = 6 or 2x - 3y - 6 = 0 here A = 2, B = -3, and C = -6

OpenStudy (raden):

wait, it like be zero as you said, haha ... |(2*6 - 3*2 - 6) / (sqrt. 2^2 + (-3)^2)| = |0/...| = 0

OpenStudy (raden):

it means the point of (6,2) is on the line 2x - 3y = 6 , so no distance of them 

OpenStudy (raden):

please recheck is the point given (6,2) or not ?

OpenStudy (anonymous):

Oh I believe it was supposed to be a trick question because the follow up question asks to explain the answer to this question. I didn't realize that (6,2) would be on the line. Thank you very much for your help!

OpenStudy (raden):

you're welcome :)